# Conditional Autoencoder for Generating BNS Waveforms with Tidal and Precession Effects

### v1

@@ Abstract
Gravitational waves from binary neutron star mergers provide insights into dense matter physics and strong-field gravity, but waveform modeling remains computationally challenging. We develop a deep generative model for gravitational waveforms from BNS mergers, covering the late inspiral, merger, and ringdown, incorporating precession and tidal effects. Using the conditional autoencoder, our model efficiently generates waveforms with high fidelity across a broad parameter space, including component masses $(m_{1},m_{2})$, spin components $(S_{1x},S_{1y},S_{1z},S_{2x},S_{2y},S_{2z})$ and tidal deformability $(\Lambda_{1},\Lambda_{2})$. Trained on $3\times 10^{5}$ waveforms from the IMRPhenomXP_NRTidalv2 waveform model, it achieves an average overlap accuracy of 99.6% on the test set. The model significantly accelerates waveform generation. For a single sample, it requires $0.12$ seconds (s), compared to $0.38$ s for IMRPhenomXP_NRTidalv2 and $0.62$ s for IMRPhenomPv2_NRTidal, making it approximately 3 to 5 times faster. When generating $10^{3}$ waveforms, the network completes the task in $0.86$ s, while traditional waveform approximation methods take over 4646464653535353 s. Our model achieves a total time of $7.48$ s to generate $10^{4}$ such waveforms, making it about 60 to 65 times faster than traditional waveform approximation methods. This speed advantage enables rapid parameter estimation and real-time gravitational wave searches. With higher precision, it will support low-latency detection and broader applications in multi-messenger astrophysics.
@@ I Introduction
Due to the significance of BNS systems, accurately modeling gravitational waveforms from their mergers is essential for both detection sensitivity and parameter estimation precision. GW searches rely on matched filtering techniques, which require highly accurate waveform templates, while extracting key physical parameters—such as masses, spins, and tidal deformabilities—demands waveform models with high fidelity. However, modeling BNS waveforms remains challenging due to complex physical effects, particularly spin precession and tidal interactions [11, 12, 13]. Waveform modeling has progressed from computationally expensive numerical relativity (NR) simulations [14, 15, 16], which solve Einstein’s equations directly, to more efficient semi-analytical methods. NR simulations provide high-precision waveforms by capturing strong-field and nonlinear effects but are too costly for large-scale parameter-space studies. Post-Newtonian (PN) approximations [17, 18] describe the inspiral phase analytically under weak-field and slow-motion assumptions but lose accuracy near merger. The effective-one-body (EOB) approach [19, 20, 21] improves upon PN by mapping the two-body problem to an effective single-body motion in a modified spacetime. With NR calibrations, EOB models achieve a balance between accuracy and efficiency. Phenomenological models (IMRPhenom) [12, 22, 23, 24] further enhance computational efficiency by fitting frequency-domain waveform templates to extensive datasets. These models enable rapid waveform generation, facilitating large-scale searches. Despite advancements, challenges remain in computational cost, accuracy, and full parameter-space coverage [25, 26, 27], limiting real-time GW detection and precise parameter estimation.
@@
The rapid development of deep learning has introduced an efficient and accurate approach to gravitational waveform modeling [28, 29, 30, 31]. With strong nonlinear fitting capabilities and high computational efficiency [32], deep learning enables high-precision waveform generation at significantly reduced computational cost. George et al. [33] first applied deep learning to BBH waveforms, achieving real-time performance beyond traditional methods. Schmidt et al. [34] used principal component analysis (PCA) with machine learning to reduce the dimensionality of EOB waveforms, improving computational efficiency. Dax et al. [35] further accelerated waveform generation by leveraging the JAX framework, enabling highly efficient computation and real-time inference. Beyond BBH systems, deep learning has shown promise in BNS and extreme mass-ratio inspirals (EMRIs) waveform modeling. Whittaker et al. [36] employed a conditional variational autoencoder (cVAE) to model post-merger signals, capturing uncertainties in the EoS for probabilistic waveform generation. Chua et al. [37] combined reduced-order modeling with deep learning to accelerate waveform generation for EMRIs, reducing the computational cost of fully relativistic waveforms for LISA data analysis by more than four orders of magnitude. These studies indicate that deep learning accelerates waveform generation and generalizes well across high-dimensional parameter spaces, providing an efficient framework for waveform generation.
@@
Despite progress in deep learning-based waveform modeling, most existing models focus on BBH systems or simplified BNS mergers, with precession and tidal effects remaining underexplored. To address this, we propose a Conditional Autoencoder (cAE) framework for rapid BNS waveform generation, with applications in GW data analysis. Our model efficiently generates waveforms conditioned on system parameters $(\Theta)$, including component masses $(m_{1},m_{2})$, spin components $(S_{1x},S_{1y},S_{1z},S_{2x},S_{2y},S_{2z})$, and tidal deformability $(\Lambda_{1},\Lambda_{2})$, while capturing the high-dimensional evolution of GW signals. Trained on a dataset of $3\times 10^{5}$ BNS waveforms from the IMRPhenomXP_NRTidalv2 [38] model, it includes both precession and tidal effects. To enhance learning efficiency, we adopt the amplitude ($A$)-phase ($\Phi$) representation, where the polarization waveforms $h_{+}(t)$ and $h_{\times}(t)$ are expressed in terms of amplitude and phase independently, to reduce data oscillation. The cAE architecture employs a dual-encoder structure, separately encoding physical parameters and waveform data, which are mapped in latent space before reconstruction. By relying solely on forward propagation, cAE achieves exceptional acceleration in large-scale waveform generation. Benchmark tests on an NVIDIA RTX 3090 GPU and an Intel Xeon Silver 4214R CPU show that for generating a single waveform, cAE requires only $0.12$ s, while IMRPhenomPv2_NRTidal and IMRPhenomXP_NRTidalv2 take $0.62$ s and $0.38$ s, respectively, making cAE approximately 5 times faster. For generating $10^{4}$ waveforms in batch, cAE completes the task in $7.48$ s, compared to $497.75$ s for IMRPhenomPv2_NRTidal and $454.51$ s for IMRPhenomXP_NRTidalv2, yielding speed-ups of about 65 and 60 times, respectively. The model’s accuracy is evaluated through waveform overlap calculations, yielding an average mismatch of $3.97\times 10^{-3}$, corresponding to accuracy 99.6%. These results demonstrate that the proposed framework enables efficient, accurate, and scalable BNS waveform generation with precession and tidal effects, making it well-suited for real-time signal detection and parameter estimation.
@@
The structure of the article is as follows: Section 2 describes the waveform representation and the construction of our dataset. Section 3 introduces the fundamental concepts of autoencoders and presents the architecture and hyperparameter settings of our neural network. Section 4 details the model training and validation process. Section 5 evaluates the accuracy and generation efficiency of our model. Finally, Section 6 provides a summary and discusses future research directions.
@@ II.1 Waveform Representation
This representation reduces data dimensionality while enhancing physical interpretability. The amplitude $A(t)$ captures the overall intensity variation of the gravitational wave, while the cumulative phase $\Phi(t)$ describes its temporal evolution, offering a depiction of the underlying dynamics. To further standardize waveform properties, we apply phase normalization,
@@ II.2 System Parameter Selection
Previous studies on gravitational waveform modeling have largely focused on BBH systems, while investigations of BNS waveforms remain relatively limited. Moreover, most deep learning models assume either non-spinning neutron stars or perfectly aligned spins, without systematically accounting for precession effects. Tidal deformation, which significantly influences the phase evolution of BNS waveforms, is often simplified using point-mass approximations, leading to the omission of crucial tidal contributions. A further limitation lies in the restricted parameter space of existing datasets. For instance, the mass range is typically constrained to $[1.2,2.0]M_{\odot}$, spin magnitudes remain small ($|S|<0.5$), and the tidal deformability parameters $\Lambda_{1},\Lambda_{2}$ have limited coverage, reducing the ability to capture variations across different EOS. These constraints hinder the generalization capability of deep learning models and limit their applicability in real GW detection. To address these challenges, we develop a machine learning framework that systematically incorporates both precession and tidal effects, enhancing waveform modeling for BNS systems.
To ensure the model effectively learns the key features of BNS waveforms, we select a set of $\Theta$ spanning a broad region of the parameter space, as summarized in TABLE 1. Specifically, we set the luminosity distance to 1 Mpc, and fix both the inclination angle and the coalescence phase to 0.
$\Theta$ Range
$m_{1}$ Uniform$[1,3]M_{\odot}$
$m_{2}$ Uniform$[1,3]M_{\odot}$
$S_{1x}$ Uniform[0.8,0.8]0.80.8[-0.8,0.8][ - 0.8 , 0.8 ]
$S_{1y}$ Uniform[0.8,0.8]0.80.8[-0.8,0.8][ - 0.8 , 0.8 ]
$S_{1z}$ Uniform[0.8,0.8]0.80.8[-0.8,0.8][ - 0.8 , 0.8 ]
$S_{2x}$ Uniform[0.8,0.8]0.80.8[-0.8,0.8][ - 0.8 , 0.8 ]
$S_{2y}$ Uniform[0.8,0.8]0.80.8[-0.8,0.8][ - 0.8 , 0.8 ]
$S_{2z}$ Uniform[0.8,0.8]0.80.8[-0.8,0.8][ - 0.8 , 0.8 ]
$\Lambda_{1}$ Uniform$[0,500]$
$\Lambda_{2}$ Uniform$[0,500]$
Table 1: Range of $\Theta$ for the BNS system in the training dataset ($m_{2}<m_{1}$).
@@ II.3 Construction of Dataset
We construct a dataset of gravitational waveforms for BNS systems to train the cAE. The dataset comprises $3\times 10^{5}$ samples generated using the IMRPhenomXP_NRTidalv2 waveform model, which incorporates both precession and tidal effects. The data generation process consists of parameter sampling, waveform computation, data preprocessing and normalization. $\Theta$ of the BNS system are drawn from uniform distributions, which is shown in TABLE 1. The time-domain GW signals $h_{+}(t)$ and $h_{\times}(t)$ are computed using pycbc.waveform.get_td_waveform[39], followed by trimming to remove leading and trailing zero values. The amplitude $A(t)$ and phase $\Phi(t)$ are then computed using Eq. (2), as shown in FIG. 1.
@@
To maintain data uniformity, waveforms are standardized and zero-padded to a fixed duration of $2$ seconds with a sampling rate of 4096 Hz, resulting in $8192$ points per sample. Each waveform segment is extracted from the final $2$ seconds before the ringdown phase, ensuring that the dataset captures the complete merger and ringdown stages while also covering part of the inspiral phase. The 2-second simulated waveform is sufficiently long to encompass the stage of prominent tidal effects preceding the merger [40], enabling the model to learn the characteristic evolution of waveforms across different dynamical regimes.
The dataset consists of three components: $X_{\text{train}}$, which contains the BNS system’s $\Theta$ as input features; $y_{A}$, representing the amplitude data; and $y_{\Phi}$, representing the phase data. During cAE training, we normalize the data to ensure stability and consistency. The $X_{\text{train}}$ are processed using Min-Max Normalization [41], which rescales the data to the range $[0,1]$, ensuring a uniform scale across different parameters and improving training stability:
@@ III Autoencoder And Conditional Autoencoder
The autoencoder (AE) is an unsupervised learning model widely used for data dimensionality reduction, feature extraction, and generative tasks. It consists of an encoder and a decoder, learning a low-dimensional representation of the data by minimizing reconstruction error. The variational autoencoder (VAE) extends this framework by introducing probabilistic modeling, enforcing a smoother latent variable distribution, which enhances the generative capability. The cAE further incorporates external conditional constraints, enabling the model to generate samples corresponding to specific data distributions based on input conditions, making it particularly relevant for GW waveform modeling.
This section first introduces the fundamental concepts of AE, VAE, and cAE, discussing their applicability to GW waveform generation. Subsequently, we provide a detailed description of the proposed cAE-based waveform generation framework, including the separate cAE architectures designed for phase and amplitude modeling, along with their respective hyperparameter settings.
@@ III.1 Concepts of Autoencoder and Conditional Autoencoder
We employ autoencoders (AEs) [43] to reduce the dimensionality of complex GW waveforms (Amplitude $A$ and Phase $\Phi$) while ensuring accurate reconstruction. As shown in FIG. 2(a), an AE consists of an input $h^{(i)}$ ($A$ or $\Phi$), an encoder $q_{\alpha}(z\mid h)$, a latent variable $z^{(i)}\in\mathbf{R}^{d}$, and a decoder $p_{\beta}({\hat{h}}\mid z)$. The encoder projects the input ($A$ or $\Phi$) into a lower-dimensional latent space $z^{(i)}$, where the dimension $d$ of $z^{(i)}$ can be adjusted based on specific task requirements. The decoder then reconstructs the $\hat{h}^{(i)}$ from the latent representation through an upsampling process. The $\alpha$ and $\beta$ refers to the learned model parameters, such as weights and biases, obtained after training. To measure the similarity between the reconstructed $\hat{h}^{(i)}$ and the target $h^{(i)}$, the AE employs the mean squared error (MSE) as the reconstruction loss:
@@
The KL divergence loss ensures that the learned latent variable distribution $q_{\alpha}(z\mid h)$ approximates a predefined prior distribution, typically a standard normal distribution $p(z)=\mathcal{N}(0,I)$:
@@
Although traditional AEs and VAEs perform well in capturing the low-dimensional structure and nonlinear features of data, their generative processes typically rely solely on the data itself. As a result, they lack the capacity to explicitly incorporate known physical priors into the latent representations. In other words, standard AE/VAE models in unsupervised learning tend to capture the dominant variations in the data, but they cannot guarantee that the generated waveforms strictly adhere to physical constraints. To address this limitation and further enhance the physical interpretability and controllability of waveform generation, we introduce the cAE [53]. In the cAE framework, additional physical parameters $\Theta$ (as shown in TABEL 1) are incorporated as conditional inputs and jointly mapped with waveform data into a low-dimensional latent space. In this manner, the cAE not only inherits the advantages of AE/VAE in nonlinear dimensionality reduction and data reconstruction, but also enables the explicit embedding of physical constraints into the latent variables, thereby generating waveforms that better reflect realistic astrophysical properties.
Our cAE employs a dual-encoder structure: one encoder processes the original $h$, while the other encodes the $\Theta$, generating latent representations $z_{A/\Phi}$ and $z_{\Theta}$, respectively. These are then mapped in the decoder to reconstruct the waveform $\hat{h}$. The total loss of training cAE consists of a reconstruction loss and a latent consistency loss:
$L_{\mathrm{cAE}}=\frac{1}{N}\sum_{i=1}^{N}\bigl{\lVert}h_{i}-\hat{h}_{i}\bigr{\rVert}_{\text{MAE}}+\lambda\frac{1}{N}\sum_{i=1}^{N}\bigl{\lVert}z_{A/\Phi,i}-z_{\Theta,i}\bigr{\rVert}_{\text{MSE}}^{2}.$ (11)
The first term measures the difference between the reconstructed waveform $\hat{h}_{i}$ and the target waveform $h_{i}$ using Mean Absolute Error (MAE), while the second term ensures consistency between the latent representations $z_{A/\Phi,i}$ and $z_{\Theta,i}$ using Mean Squared Error (MSE). The hyperparameter $\lambda$ balances the two losses, in our study, we fixed it as 1. In the following, we provide a detailed description of our model architecture.

### v4

# Conditional Autoencoder for Generating Binary Neutron Star Waveforms with Tidal and Precession Effects

@@ Abstract
Gravitational waves from binary neutron star mergers provide insights into dense matter physics and strong-field gravity, but waveform modeling remains computationally challenging. We develop a deep generative model for gravitational waveforms from binary neutron star mergers, covering the late inspiral, merger, and ringdown, incorporating precession and tidal effects. Using the conditional autoencoder, our model efficiently generates waveforms with high precision across a broad parameter space, including component masses $(m_{1},m_{2})$, spin components $(S_{1x},S_{1y},S_{1z},S_{2x},S_{2y},S_{2z})$ and tidal deformability $(\Lambda_{1},\Lambda_{2})$. Trained on $1\times 10^{6}$ waveforms from the IMRPhenomXP_NRTidalv2 waveform model, our model achieves a mean mismatch of $2.13\times 10^{-3}$. The model accelerates waveform generation. For a single sample, it requires $0.12$ seconds (s), compared to $0.66$ s for IMRPhenomXP_NRTidalv2 making it approximately 5 times faster. When generating 1000 waveforms, the network completes the task in $0.75$ s, while IMRPhenomXP_NRTidalv2 requires $7.12$ s, making it approximately 10 times faster. This speed advantage enables rapid parameter estimation and real-time gravitational wave searches. With higher precision, it will support low-latency detection and broader applications in multi-messenger astrophysics.
@@ I Introduction
Due to the significance of BNS systems, accurately modeling gravitational waveforms from their mergers is essential for both detection sensitivity and parameter estimation precision. GW searches rely on matched filtering techniques, which require highly accurate waveform templates, while extracting key physical parameters—such as masses, spins, and tidal deformabilities—demands waveform models with high precision. However, modeling BNS waveforms remains challenging due to complex physical effects, particularly spin precession and tidal interactions [11, 12, 13]. Waveform modeling has progressed from computationally expensive numerical relativity (NR) simulations [14, 15, 16], which solve Einstein’s equations directly, to more efficient semi-analytical methods. NR simulations yield high-precision waveforms by capturing strong-field, nonlinear effects but are too costly for large-scale parameter-space studies. Post-Newtonian (PN) approximations [17, 18] describe the inspiral phase under weak-field, slow-motion assumptions but lose accuracy near merger. The effective-one-body (EOB) approach [19, 20, 21] improves upon PN by mapping the two-body problem to a single-body motion in a modified spacetime, and with NR calibration, balances accuracy and efficiency. Phenomenological models (IMRPhenom) [12, 22, 23, 24] further enhance efficiency by fitting frequency-domain templates to extensive datasets, enabling rapid waveform generation and facilitating large-scale searches. Despite advancements, challenges remain in computational cost, accuracy, and full parameter-space coverage [25, 26, 27], limiting real-time GW detection and precise parameter estimation.
@@
The rapid development of deep learning has introduced an efficient and accurate approach to gravitational waveform modeling [28, 29, 30, 31]. With strong nonlinear fitting capabilities and high computational efficiency [32], deep learning enables high-precision waveform generation at significantly reduced cost. George et al. [33] first applied deep learning to BBH waveforms and achieved real-time performance beyond traditional methods. Schmidt et al. [34] used PCA with machine learning to reduce the dimensionality of EOB waveforms and improve efficiency. Dax et al. [35] accelerated waveform generation using the JAX framework for highly efficient computation and real-time inference. Beyond BBH systems, deep learning has been applied to BNS and EMRI waveform modeling. Whittaker et al. [36] used a cVAE to model post-merger signals with EoS uncertainties. Chua et al. [37] combined reduced-order modeling with deep learning to accelerate EMRI waveform generation, reducing the cost by over four orders of magnitude. These works show that deep learning accelerates waveform generation and handles high-dimensional parameter spaces effectively.
@@
Despite progress in deep learning-based waveform modeling, most existing models focus on BBH systems or simplified BNS mergers, with precession and tidal effects remaining underexplored. To address this, we propose a Conditional Autoencoder (cAE) model for rapid BNS waveform generation, with applications in GW data analysis. Our model efficiently generates waveforms conditioned on system parameters $(\Theta)$, including component masses $(m_{1},m_{2})$, spin components $(S_{1x},S_{1y},S_{1z},S_{2x},S_{2y},S_{2z})$, and tidal deformability $(\Lambda_{1},\Lambda_{2})$, while capturing the high-dimensional evolution of GW signals. Trained on a dataset of $3\times 10^{5}$ BNS waveforms from the IMRPhenomXP_NRTidalv2 [38] model, it incorporates both precession and tidal effects. To enhance learning efficiency, we adopt the amplitude ($A$)-phase ($\Phi$) representation, where $h_{+}(t)$ and $h_{\times}(t)$ are expressed in terms of amplitude and phase independently, to reduce data oscillation. The cAE architecture employs a dual-encoder structure, separately encoding physical parameters and waveform data, which are mapped in latent space before reconstruction. By relying solely on forward propagation, cAE achieves high acceleration in large-scale waveform generation. Benchmark tests show that generating a single waveform takes 0.12 s with cAE on an NVIDIA A800 80 GB GPU, compared to 0.66 s with IMRPhenomXP_NRTidalv2 on two Intel Xeon Silver 4214R CPUs (24 cores), corresponding to a speed-up of about 5 times; for a batch of 10³ waveforms, cAE needs 0.75 s compared with 7.12 s, yielding roughly 10 times speed-up. The model’s accuracy is evaluated through waveform overlap calculations, yielding an average mismatch $2.13\times 10^{-3}$, corresponding to accuracy 99.79%. These results demonstrate that the proposed model enables efficient, accurate, and scalable BNS waveform generation with precession and tidal effects, making it well-suited for real-time signal detection and parameter estimation.
@@
The structure of the article is as follows: Section II describes the waveform representation and the construction of our dataset. Section III introduces the fundamental concepts of autoencoders and presents the architecture and hyperparameter settings of our neural network. Section IV details the model training and validation process. Section V evaluates the accuracy and generation efficiency of our model. Finally, Section VI provides a summary and discusses future research directions.
@@ II.1 Waveform Representation
This representation reduces data Oscillatory while enhancing physical interpretability. The amplitude $A(t)$ captures the overall intensity variation of the gravitational wave, while the cumulative phase $\Phi(t)$ describes its temporal evolution, offering a depiction of the underlying dynamics. To further standardize waveform properties, we apply phase normalization,
@@ II.2 System Parameter Selection
Previous studies on gravitational waveform modeling have primarily focused on BBH systems, while investigations of BNS waveforms remain relatively limited. Most existing deep learning models assume binary neutron stars with spins aligned to the orbital angular momentum, and thus do not systematically account for spin precession effects. Additionally, although tidal deformation can significantly affect the phase evolution of BNS waveforms, it is often simplified using point-mass approximations, which may result in the omission of tidal contributions. In this work, we fix the luminosity distance to 1 Mpc, and set both the inclination angle and the coalescence phase to zero. The remaining parameters are listed in TABLE 1, where the spin magnitude ranges are chosen based on the benchmark settings provided in [39].
Table 1: Range of the sampling parameters $\Theta$ for the BNS training set (with $m_{2}<m_{1}$).
   $\Theta$    Description    Range
[…]
   $|\vec{s}_{1}|,\;|\vec{s}_{2}|$    Spin magnitudes of two neutron stars    Uniform[0, 0.5][0,\,0.5]
   $\vec{s}_{1}/\left|\vec{s}_{1}\right|,\vec{s}_{2}/\left|\vec{s}_{2}\right|$    Spin directions (unit vectors)    Isotropic over 3D sphere
   $\Lambda_{1},\;\Lambda_{2}$    Tidal deformabilities of two neutron stars    Uniform[0, 500][0,\,500]
@@ II.3 Construction of Dataset
We construct a dataset of gravitational waveforms for BNS systems to train and evaluate our conditional autoencoder model. The training set consists of $1\times 10^{6}$ samples, while the test set comprises $1\times 10^{5}$ samples. Both sets are generated using the IMRPhenomXP_NRTidalv2 waveform model [39], which incorporates precession and tidal effects.
The data generation process includes parameter sampling, waveform computation, preprocessing, and normalization. The sampled parameters $\Theta$ include component masses, spin vectors, and tidal deformabilities, with ranges summarized in TABLE 1. Component masses $m_{1}$ and $m_{2}$ ($m_{2}<m_{1}$) are independently drawn from a uniform distribution over $1,\,3]\,M_{\odot}$, with values reordered post-sampling to ensure $m_{1}>m_{2}$. The dimensionless spin magnitudes $|\vec{s}_{1}|$ and $|\vec{s}_{2}|$ are sampled uniformly in [0, 0.5][0,\,0.5], and their directions are drawn from an isotropic distribution on the unit sphere to allow for generic spin precession. Tidal deformability parameters $\Lambda_{1}$ and $\Lambda_{2}$ are sampled uniformly in [0, 500][0,\,500].
Both training and test sets are constructed using the same stochastic (non-grid) sampling strategy without predefined step sizes. This randomized approach avoids artifacts introduced by grid-based sampling, enhances waveform diversity, and promotes broad coverage of the high-dimensional parameter space—factors essential for improving generalization and avoiding overfitting in deep learning models. The distributions of sampled parameters in both the training and testing sets are visualized in Appendix A, FIGs. 14 and 15, respectively.
The time-domain GW signals $h_{+}(t)$ and $h_{\times}(t)$ are computed using pycbc.waveform.get_td_waveform [40], followed by trimming to remove leading and trailing zero values. Each waveform is standardized to a fixed duration of $2$ seconds with a sampling rate of 4096 Hz. This segment is taken from the final two seconds before the end of the ringdown, ensuring that the dataset captures the complete merger and ringdown stages while also covering part of the inspiral. The chosen duration is sufficiently long to encompass the stage of prominent tidal effects preceding the merger [41], enabling the model to learn the characteristic evolution of waveforms across different dynamical regimes.
@@
To maintain data consistency, waveforms are to a fixed duration of $2$ seconds with a sampling rate of 4096 Hz. Each waveform segment is taken from the 2 seconds before the end of the ringdown, ensuring that the dataset captures the complete merger and ringdown stages while also covering part of the inspiral. The 2-second simulated waveform is sufficiently long to encompass the stage of tidal effects preceding the merger [41].
The dataset consists of three components: $X_{\text{train}}$, which contains the BNS system’s $\Theta$ as conditional inputs; $y_{A}$, representing the amplitude data; and $y_{\Phi}$, representing the phase data. During cAE training, we normalize the data to ensure stability and consistency. The $X_{\text{train}}$ are processed using Min-Max Normalization [42], which rescales the data to the range $[0,1]$, ensuring a uniform scale across different parameters and improving training stability:
@@ III Autoencoder And Conditional Autoencoder
The autoencoder (AE) is an unsupervised learning model widely used for data dimensionality reduction, feature extraction, and generative tasks. It consists of an encoder and a decoder, learning a low-dimensional representation of the data by minimizing reconstruction error. The variational autoencoder (VAE) extends this model by introducing probabilistic modeling, enforcing a smoother latent variable distribution, which enhances the generative capability. The cAE further incorporates external conditional constraints, enabling the model to generate samples corresponding to specific data distributions based on input conditions, making it particularly relevant for GW waveform modeling.
This section first introduces the fundamental concepts of AE, VAE, and cAE, discussing their applicability to GW waveform generation. Subsequently, we provide a detailed description of the proposed cAE-based waveform generation model, including the separate cAE architectures designed for phase and amplitude modeling, along with their respective hyperparameter settings.
@@ III.1 Concepts of Autoencoder and Conditional Autoencoder
We employ autoencoders (AEs) [44] to reduce the dimensionality of complex GW waveforms (Amplitude $A$ and Phase $\Phi$) while ensuring accurate reconstruction. As shown in FIG. 2(a), an AE consists of an input $h^{(i)}$ (AA or $\Phi$), an encoder $q_{\alpha}(z\mid h)$, a latent variable $z^{(i)}\in\mathbf{R}^{d}$, and a decoder $p_{\beta}({\hat{h}}\mid z)$. The encoder projects the input ($A$ or $\Phi$) into a lower-dimensional latent space $z^{(i)}$, where the dimension $d$ of $z^{(i)}$ can be adjusted based on specific task requirements. The decoder then reconstructs the $\hat{h}^{(i)}$ from the latent representation through an upsampling process. The $\alpha$ and $\beta$ refers to the learned model parameters, such as weights and biases, obtained after training. To measure the similarity between the reconstructed $\hat{h}^{(i)}$ and the target $h^{(i)}$, the AE employs the mean squared error (MSE) as the reconstruction loss:
@@
The KL divergence loss ensures that the learned latent variable distribution $q_{\alpha}(z\mid h)$ approximates a predefined prior distribution, typically a standard normal distribution $p(z)=\mathcal{N}(0,I)$ where $I$ is the identity matrix:
@@
Although traditional AEs and VAEs perform well in capturing the low-dimensional structure and nonlinear features of data, their generative processes typically rely solely on the data itself. As a result, they lack the capacity to explicitly incorporate known physical priors into the latent representations. In other words, standard AE/VAE models in unsupervised learning tend to capture the dominant variations in the data, but they cannot guarantee that the generated waveforms strictly adhere to physical constraints. To address this limitation and further enhance the physical interpretability and controllability of waveform generation, we use the cAE [54]. In the cAE model, additional physical parameters $\Theta$ (as shown in TABLE 1) are incorporated as conditional inputs and jointly mapped with waveform data into a low-dimensional latent space. In this manner, the cAE not only inherits the advantages of AE/VAE in nonlinear dimensionality reduction and data reconstruction, but also enables the explicit embedding of physical constraints into the latent variables, thereby generating waveforms that better reflect realistic astrophysical properties.
In the following, we provide a detailed description of our cAE model architecture.